OpenStudy (anonymous):

5/4x+1/8x=5/8+x ?? I am so bad at fractions

OpenStudy (anonymous):

like this?: \[\frac{5}{4}x+\frac{1}{8}x=\frac{5}{8}+x\]

OpenStudy (anonymous):

yes but how do you solve it??

OpenStudy (anonymous):

ok, multiply everything by 8: 10x+x=5+8x 3x=5 x=5/3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

why do we multiply itby 8 b/c it is the common denominator?

OpenStudy (anonymous):

its the least common multiple of the denominators, so it gets rid of all the fractions.

OpenStudy (anonymous):

ok I will kepp that in mind thank you..

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

do you k now anything about the elimination method?

OpenStudy (anonymous):

do you have a specific problem?

OpenStudy (anonymous):

just getting it to opposite and yes! the 5x+6y=2 10x+12y=4

OpenStudy (anonymous):

well those equations are multiples of each other

OpenStudy (anonymous):

so elimination isnt really useful here

OpenStudy (anonymous):

that answer calls for use of the elimination

OpenStudy (anonymous):

to eliminate either x or y you multiply the first equation by -2, then add. But then you get 0=0

OpenStudy (anonymous):

there are actually many solutions to that system

OpenStudy (anonymous):

theyre both the same equation

OpenStudy (anonymous):

there is an option for no solutions or many solutions

OpenStudy (anonymous):

so many solutions woiuld fit

OpenStudy (anonymous):

there are many solutions, because theyre both the same equation so you can pick any x and find a corresponding y to make the equation true

OpenStudy (anonymous):

thank you again

OpenStudy (anonymous):

but something like 2x+3y=15 4x+y=20 you would multiply the first equation by -2 and then add them together to eliminate x -4x-6y=-30 4x+y=20 -5y=-10 y=2 then plug back in to find x 2x+6=15 x=9/2

OpenStudy (anonymous):

then the solution is unique